Stuart Manufacturing Company was started on January 1, 2018, when it acquired $78,000 cash by issuing common stock. Stuart immediately purchased office furniture and manufacturing equipment costing $9,100 and $26,900, respectively. The office furniture had an eight-year useful life and a zero salvage value. The manufacturing equipment had a $3,200 salvage value and an expected useful life of three years. The company paid $11,800 for salaries of administrative personnel and $15,200 for wages to production personnel. Finally, the company paid $15,120 for raw materials that were used to make inventory. All inventory was started and completed during the year. Stuart completed production on 4,900 units of product and sold 3,940 units at a price of $15 each in 2018. (Assume that all transactions are cash transactions and that product costs are computed in accordance with GAAP.)

Required

  1. Determine the total product cost and the average cost per unit of the inventory produced in 2018. (Round "Average cost per unit" to 2 decimal places.)

  2. Determine the amount of cost of goods sold that would appear on the 2018 income statement. (Do not round intermediate calculations.)

  3. Determine the amount of the ending inventory balance that would appear on the December 31, 2018, balance sheet. (Do not round intermediate calculations.)

  4. Determine the amount of net income that would appear on the 2018 income statement. (Round your answer to the nearest dollar amount.)

  5. Determine the amount of retained earnings that would appear on the December 31, 2018, balance sheet. (Round your answer to the nearest dollar amount.)

  6. Determine the amount of total assets that would appear on the December 31, 2018, balance sheet. (Round your answer to the nearest dollar amount.)
Answer #1

1. Total product cost = wages to production personnel+ raw materials that were used to make inventory + Depreciation on equipment = 15200+ 15,120 +7900 = 38220

Average cost per unit = 38220/ 4900 = 7.8 per unit

2. Cost of goods sold = 7.8 * 3940 =  30732

3. Ending inventory = 38220 - 30732 = 7488

4. Net Income = 15430

Sales 59100
Less: Cost of goods sold 30732
Gross Margin 28368
Less: salaries of administrative personnel 11800
Depreciation expense - office furniture 1138
Net Income 15430

6. Total assets = 93430

Cash
[78000-9100-26900-11800-15200-15120+59100]
58980
Equipment 19000
Office furniture 7962
Ending inventory 7488
Total assets 93430